OpenStudy (anonymous):

prove: csc^2 X-1/ csc^2 X = cos^2 X

OpenStudy (mathmath333):

is it \(\large\tt \begin{align} \color{black}{ \csc^2 x-\dfrac{1}{\csc^2 x}=\cos^2 x \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

csc^2x-1 = cot^2x or coot^2x/csc^2x or (cos^2x/sin^2x)*(1/csc^2x) or (cos^2x/sin^2x)*sin^2x =cos^2x